OpenStudy (anonymous):

what is the factor of 2x4 + 22x3 + 60x2?

OpenStudy (anonymous):

\[\huge 2x^4 + 22x^3 + 60x^2\] from the given expression we can see that \[\huge 2x^2 \] is common in all the terms hence we take it common. \[\huge 2x^4 + 22x^3 + 60x^2 = 2x^2(x^2 + 11x + 30)\] \[\huge = 2x^2(x^2 + 6x+5x + 30) \] \[\huge = 2x^2[x(x + 6)+5(x + 6)]\] \[\huge = 2x^2(x+5)(x + 6)\] are the required factors

OpenStudy (anonymous):

@calidreamer213

OpenStudy (anonymous):

my choices were 2x3, x +5, x+4, x3

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

The factors what i have obtained are correct one

OpenStudy (anonymous):

Well your factors wasn't on the list of correct answers

OpenStudy (anonymous):

@dpasingh